# Satellite Federated Fine-Tuning for Foundation Models in Space Computing Power Networks

### v2

@@ VI-A1 LEO Constellation Configurations
We conduct simulations using an $80/4/1$ Walker constellation setup. This configuration consists of 4 orbits, with 20 satellites evenly distributed in each orbit. All satellites operate at an altitude of 550 km. The parameters for the transmitters and receivers, as well as the constellation, are shown in Table I.
@@ VI-A3 Baselines
  • Training on Satellites. Each satellite trains a local model on the collected dataset, and their head weights are aggregated through ISLs.
  • Training on the Ground. Download satellite data to the ground and train the model on a ground server. Note that the results of the simulations do not account for the transmission time of raw data.
TABLE I: Simulation parameters
@@ VI-B1 Training Time
The training time of the proposed framework is compared with two baselines as shown in Fig 8. It can be observed that the proposed framework achieves approximately $33\%$ reduction in training time when compared to the fine-tuning approach on satellites. This reduction in training time can be attributed to the fact that, in the proposed framework, the majority of the computation workload is handled by the PS, which operates at a much higher speed than the computing devices of the satellites. This highlights that when the model training cost is substantial, the time required for satellites to train the entire model locally often exceeds the time it takes to transmit training data between the GSs and the satellites. This discrepancy is primarily due to the limited processing power of the satellites, which makes local training less efficient. In contrast, by offloading the heavy computation tasks to the PS and only transmitting small training information, the proposed framework reduces the overall training time, thus optimizing the overall performance of the system.
VI-B2 Model Size
@@
Figure 9: Average training time versus model size.
VI-B3 The Overhead of Satellite-Ground Transmission
@@
    Bridge dataset [39]. This dataset is for object detection and consists of 500 images, each containing at least one bridge from different regions worldwide. All images are $4,800\times 2,843$ pixels.
@@
    Aerial image segmentation dataset [40]. This is used in image segmentation, where a semantic class label is assigned to each pixel as a basis for automatic map generation. Each image is approximately $3000\times 3000$ pixels.
@@
    OSCD [41]. It contains 24 pairs of multispectral images taken between 2015 and 2018 and provides pairs of satellite images in 13 bands that can be used for change detection. The spatial resolution of the image varies between 10m, 20m, and 60m. The image size is $600\times 600$ pixels.
@@
We present the results in Fig. 10. For datasets with 13 bands, we convert them into images with 3 bands of the same size to display them in the figure. The figure displays the ratio of the transmitted data size to the original data size when the original images vary in size. This ratio provides a clear indication of how much data is being transmitted between the satellite and the GS relative to the size of the original dataset. As the size of the original images increases, this ratio decreases, which can be attributed to the information extraction capability of the FMs. Several points in the figure are marked, each representing a specific scenario where the data transmission ratio is calculated for a given dataset. For example, in the simulations conducted on the SegMunich dataset, the transmitted data size is approximately $2\%$ of the size of the original dataset.
@@ VI-B4 Communication Strategy
    Strategy 2: Satellite-ground transmission without inter-satellite cooperation [17]. Each satellite only transmits local data until the PS receives all data.
@@
Figure 11: Learning performance versus training time in different strategies. Strategy 1: sequential intra-orbit transmission. Strategy 2: Satellite-ground transmission without inter-satellite cooperation. Strategy 3: Inter-orbit gossip transmission.

### v3

@@ VI-A1 LEO Constellation Configurations
We conduct simulations under 80/4/180/4/180 / 4 / 1, 160/4/1160/4/1160 / 4 / 1 and 200/10/1200/10/1200 / 10 / 1 Walker constellation setup. 80/4/180/4/180 / 4 / 1 constellation consists of 4 orbits, with 20 satellites evenly distributed in each orbit. 160/4/1160/4/1160 / 4 / 1 constellation is composed of 4 orbits with 40 satellites in each orbit. 200/10/1200/10/1200 / 10 / 1 constellation contains 10 orbits, with 20 satellites in each orbit. All satellites operate at an altitude of 590 km. The parameters for the transmitters and receivers, as well as the constellation, are shown in Table II.
@@ VI-A3 Baselines
  • Training on satellites(ToS). This represents a traditional satellite-only federated learning paradigm where each satellite independently trains a local model using its on-board data, and task-specific head parameters are aggregated through the proposed communication strategy. It is chosen to highlight the inefficiency of on-board fine-tuning for large foundation models due to constrained computational resources. This baseline directly contrasts with our framework’s satellite-ground collaborative design, emphasizing the value of offloading heavy computations to ground servers.
  • Centralized training on the ground. This baseline simulates a centralized training scheme where raw satellite data is downloaded to a ground server for model training, with no on-board processing. It is included to compare the accuracy of centralized training and federated training.
  • DSFL [24]. A decentralized framework where satellites collaborate via intra-orbit and inter-orbit ISLs without a central ground server.
  • Satellite cluster FL [22].The on-Board federated learning for satellite clusters with inter-Satellite links algorithm is a centralized framework, which uses intra-orbit links to construct a stable loop topology for low-orbit satellite constellations, and allows satellites in the same orbit to accumulate local gradients and neighbor gradients through an incremental aggregation mechanism and transmit them to the parameter server by sink nodes, and selects the appropriate sink nodes in the communication window based on predictive routing.
TABLE II: Simulation parameters
@@
$(k,\alpha)$ $(0.0751dB/km\cdot(mm/h)^{(-\alpha)},0.083)$ Rain attenuation coefficient
@@ VI-B1 Training Time
The training time of the proposed framework is compared with baselines as shown in Fig 8. It can be observed that the proposed framework reduces training time to 33% of on-board training time. This reduction in training time can be attributed to the fact that, in the proposed framework, the majority of the computation workload is handled by the PS, which operates at a much higher speed than the computing devices of the satellites. This highlights that when the model training cost is substantial, the time required for satellites to train the entire model locally often exceeds the time it takes to transmit training data between the GSs and the satellites. This discrepancy is primarily due to the limited processing power of the satellites, which makes local training less efficient. In contrast, by offloading the heavy computation tasks to the PS and only transmitting small training information, the proposed framework reduces the overall training time, thus optimizing the overall performance of the system.
The latency of traditional satellite federated learning is mainly dominated by computational latency, with transmission latency accounting for a very small proportion. The computing speed of satellites is much lower than that of ground servers, while the backbone network of remote sensing models requires a lot of computing power. When training on a satellite, it takes a lot of time to complete forward propagation and back propagation. The amount of updated model parameters is very small (for example, the task header parameters are only a few MB) and the transmission rate of ISLs can reach Gbps. Therefore, the impact of transmission delay on the total delay can be ignored. Therefore, the total latency in different on-board federated learning algorithms is roughly the same.
VI-B2 Time Components
Training time can be classified into five types. Their specific definitions, characteristics, and functions are as follows:
  • On-board computation: It refers to the computing tasks executed locally on the satellite, mainly involving the computation of the embedding layer and task-specific heads.
  • Terrestrial computing: It refers to the computing tasks executed by PS, such as the computation of the model’s backbone network.
  • Intra-Orbit transmission: Intraorbit transmission refers to data exchange among satellites within the same orbit, which functions to aggregate embedding vectors and synchronize task-specific head parameters. Within this transmission procedure, embedding vectors constitute the predominant payload, with their volume substantially exceeding that of head parameters. This discrepancy arises primarily from two key factors: first, the size of individual embedding vectors, measured in megabytes is significantly larger than that of head parameters, which are on the order of kilobytes; second, the quantity of embedding vectors is proportional to the number of data samples processed on each satellite, whereas each satellite maintains only a single task-specific head.
  • Inter-orbit Transmission: It refers to the data transmission between satellites in different orbits, which is used for the aggregation and broadcasting of global task header parameters.
  • Satellite-ground transmission: It refers to the data transmission between the satellite and the GS, which is used for downloading the embedding vectors and uploading the feature vectors.
VI-B3 The Effect of Data Volume
We analyzed the impact of the local dataset size of each satellite on time components as shown in Fig. 9. On-board computation mainly involves the computation of the embedding layer and task-specific heads. Its time increases linearly with the growth of sample size. This is because satellites need to perform embedding layer and loss calculations for each sample. However, the on-board computation time accounts for a small proportion due to the relatively small on-board computing tasks. The data volume also has an impact on terrestrial computing, but ground servers can alleviate the time increase caused by data volume by dynamically allocating server clusters. Intra-orbit transmission refers to data transmission between satellites within the same orbit, used for embedding vector aggregation and task header parameter synchronization. The time for embedding vector broadcasting increases linearly with the increase in sample size. This is because satellites in the same orbit need to exchange the generated embedding vectors, and even with a parallel strategy, the total transmission time still increases with the sample size. Inter-orbit transmission is less affected by the sample size, as it is used for global task head parameter aggregation. Satellite-ground transmission is used for downloading embedding vectors and uploading feature vectors. It is affected linearly by the number of samples. In summary, the time of on-board computing, intra-orbit transmission, and satellite-ground transmission increases linearly with data volume. Terrestrial computing and inter-orbit transmission are less affected by data volume. Traditional methods exhibits a longer training time and the training time in traditional methods increases linearly with data volume.
Figure 9: The effect of data volume
VI-B4 The Effect of Data Rate
The SGL data rate and ISL data rate mainly affect the total training time by influencing intra-orbit transmission, inter-orbit transmission, and satellite-ground transmission, but have no direct impact on onboard computing and terrestrial computing.
The SGL data rate directly determines the transmission efficiency between the satellite and the ground station, mainly affecting satellite-ground transmission, and having almost no impact on other time components. The satellite-ground transmission time taken is negatively correlated with the SGL data rate when the total data volume of embedding vectors and feature vectors is fixed. The Fig. 10 shows that regardless of the ISL rate being 10 Gbps, 1 Gbps, or 500 Mbps, the transmission time significantly decreases as the SGL data rate increases from 40 to 200 Mbps.
Figure 10: The effect of data rate
The ISL data rate influences the efficiency of data exchange between satellites, mainly affecting intra-orbit and inter-orbit transmission, and has no direct impact on satellite-to-ground transmission or computing. Intra-orbit transmission achieves embedded vector broadcasting and task header parameter aggregation through ISLs, using the Ring Allreduce parallel strategy. The time is negatively correlated with the ISL rate. Fig. 10 shows that at the same SGL rate, the transmission time for a 10 Gbps ISL is shorter than that for a 1 Gbps and 500 Mbps ISL. Inter-orbit transmission achieves global task header parameter aggregation through inter-satellite links (ISLs), and its time is also negatively correlated with the ISL rate. When ISL data rate reach 10Gbps, satellite-ground transmission time accounts for a huge proportion of total transmission time.
VI-B5 The Effect of Constellation Size
When there are different number of satellites in each orbit, An increase in the number of satellites per orbit will increase the transmission time for embedding vector broadcasting, embedding vector downloading and task header parameter aggregation. Even with a parallel transmission strategy, the intra-orbit transmission time will rise due to the increased number of embedding vectors that need to be exchanged among satellites in the same orbit. In Fig. 11, as the number of satellites per orbit increases, the transmission time for embedding vector broadcasting shows a linear growth. An increase in the number of satellites per orbit also leads to an increase in the total data volume for satellite-ground transmission. Because more satellites lead to more connection chances, satellite-ground transmission time does not increase linearly.
An increase in the number of orbits affects the path selection of inter-orbit data transmission. When performing global task header parameter aggregation, the transmission delay is increased linearly. More orbits mean that more satellites need to communicate with the ground station, increasing the overhead of SGLs. The increase of satellite-ground transmission time is not linear because more satellites lead to more connection windows.
VI-B6 Model Size
@@
Figure 11: The effect of constellation size
Figure 12: Total training time versus model size.
VI-B7 The Overhead of Satellite-Ground Transmission
@@
    Bridge dataset [42]. This dataset is for object detection and consists of 500 images, each containing at least one bridge from different regions worldwide. All images are 4,800×2,8434,800$\times$ 2,8434 , 800 × 2 , 843 pixels.
@@
    Aerial image segmentation dataset [43]. This is used in image segmentation, where a semantic class label is assigned to each pixel as a basis for automatic map generation. Each image is approximately $\times 3000$ 3000 pixels.
@@
    OSCD [44]. It contains 24 pairs of multispectral images taken between 2015 and 2018 and provides pairs of satellite images in 13 bands that can be used for change detection. The spatial resolution of the image varies between 10m, 20m, and 60m. The image size is $\times 600$ 600 pixels.
@@ VI-B8 Communication Strategy
    Strategy 2: Satellite-ground transmission without inter-satellite cooperation(Satellite-ground transmission without ISC) [18]. Each satellite only transmits local data until the PS receives all data.
@@
Figure 14: Learning performance versus training time in different strategies.
VI-B9 Link Utilization
The satellite-ground communication network is modeled as a directed acyclic graph with capacity constraints, and the maximum flow algorithm is used to allocate transmission tasks. In each time slice, all available SGLs are used in parallel and fully fill the communication window. And the embedding vectors aggregated in the orbit are transmitted simultaneously for satellites. For example, after satellites in a certain orbit aggregate data through a ring broadcast, it can be transmitted by any satellite in the communication window, without the need for other satellites to wait, thus reducing the link idle time and improve the link utilization.
During intra-orbit transmission, the data is divided into N blocks. In the scatter-reduce phase, each satellite sends data to its neighbors and receives data from its neighbors, completing local aggregation synchronously. In the collection phase, the aggregated data blocks are transmitted in reverse. During the entire process, all links are occupied at the same time, and there are no idle links, and the link utilization rate is close to 100%.